Scott wants to make 15 L of a 68% acid solution by mixing together a 75% acid solution and a 60% acid solution. How much of each solution must be used?
Answer by ptaylor(2198)

Let x=amount of 75% solution used
Then 15-x=amount of 60% solution used
The amount of pure acid in the 75% solution (0.75x) plus the amount of pure acid in the 60% solution (0.60(15-x)) has to equal the amount of pure acid in the final mixture (0.68*15), so our equation to solve is:
0.75x+0.60(15-x)=0.68*15 get rid of parens
0.75x+9-0.60x=10.2 subtract 9 from each side
0.75x+9-9-0.60x=10.2-9 colect like terms
0.15x=1.2 divide each side by 0.15
x=8 L------------------------amount of 75% solution used
15-x=15-8=7 L-----------------amount of 60% solution used
CK
0.75*8+0.60*7=0.68*15
6+4.2=10.2
10.2=10.2
Hope this helps----ptaylor